OpenStudy (anonymous):

Find the radian measure of the central angle of the circle of radius 6 centimeters that intercepts an arc of length 32 centimeters.

OpenStudy (ash2326):

let the radian angle be x then \[x\times r=\text{arc length}\] r=6 cm arc length=32 find x in radians?

OpenStudy (anonymous):

32*6=192 degrees

OpenStudy (anonymous):

OK so I can convert this?

OpenStudy (anonymous):

192 degrees (pi/180)?

OpenStudy (anonymous):

Do you see errors? yet?

OpenStudy (ash2326):

\[x\times 6=32\] \[x=\frac {32}{6}\] and you don't have to convert to radians, it's already in radians when x=2pi then the arc length is perimeter

OpenStudy (anonymous):

That's less complicated than expected. I see where I confused myself now.

OpenStudy (ash2326):

good:)

OpenStudy (anonymous):

My solution sheet has theta =something else

OpenStudy (anonymous):

:( uh oh

OpenStudy (ash2326):

what's the answer given as?

OpenStudy (anonymous):

theta =16/3

OpenStudy (anonymous):

Recall the formula s = rθ, where θ is measured in radians, s is the length of the arc of the circle intercepted by the central angle θ, and r is the radius. ( I think this will help).

OpenStudy (anonymous):

Any body wanna comment?